JJ Corporation purchased a machine on July 1, 20x0 for $750,000. The machine was estimated to have a useful life of 10 years with a salvage value of $46,000. JJ uses the sum-of-years-digits method to depreciate this class of equipment. During 20x3, it became apparent that the machine would become uneconomical to operate after December 31, 20x6, and that the machine would have no salvage value. What amount should be reported for Depreciation Expense as of December 31, 20x3 for this machine?
Answer #1

Sum of the year digit = 10+9+8+7+6+5+4+3+2+1 = 55

Accumulated depreciation (July 1 20X0 to June 30 20x2) = (750000-46000)*19/55 = 243200

June 30 20x2 to Dec 31 20x2 = (750000-46000)*8/55*6/12 = 51200

Total Accumulated depreciation = 243200+51200 = 294400

Depreciation expense as of December 31,20X6 = (750000-294400)*4/10 = 182240